Yellow light emitted from a sodium lamp has a wavelength (.) of 580 nm. Calculate frequency (.) and wave number of the yellow light?

Grade:12

3 Answers

Sunil Kumar FP
9 years ago
Yellow light emitted from a sodium lamp has a wavelength (.) of 580 nm
wavelength=580nm =580*10^-9m
we have
frequency =velocity of light /wavelength
=3*10^8/580*10^-9
=5.17*10^14 per second
wavenumber =1/wavelegth
=1/580*10^-9
=1.724*10^6 m^-1
swarup
5 years ago
V=c÷gama 
:v=3×10^8÷580×10^-9
:v=5.17×10^14s^-1
:wave no. of yellow light=v=1÷gama
                                               =1÷580×10^-9
                                               =1.72×10^6
Kushagra Madhukar
3 years ago
Dear student,
Please find the solution to your problem below.
Yellow light emitted from a sodium lamp has a wavelength (.) of 580 nm wavelength=580nm =580*10-9 m
we have frequency = velocity of light /wavelength = 3 x 108/580 x 10-9 = 5.17 x 1014 per second
wavenumber = 1/wavelegth =1/580x10-9 =1.724 x 106 m-1

Hope it helps.
Thanks and regards,
Kushagra